an angle measures 40 degrees more than the measure of a complementary angle.what is the measure of the greater angle?
Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

The sum of the two complimentary angles is 90 deg
let the angle be x
The other angle will be x+40
x+x+40 = 90
2x+40=90
2x=90-40
2x=50
x=50/2
x=25

the bigger angle is greater by 40 deg.
25+40 = 65 degrees.

Answer by rfer(16322)

x+x+40=90
2x+40=90
2x=50
x=25 degrees
x+40=65 degrees